A screen is placed 49.0 cm from a single slit, which is illuminated with light of wavelength 684 nm. If the distance between the first and third minima in the diffraction pattern is 2.80 mm, what is the width of the slit?
Answer #1

λ = 684 nm, D = .49 m

You need to use the formula -

λ = ws/D

Rearrange to get in terms of s(slit spacing)

s = λD/w

Throw in your values (remember to convert to m) -

s= [2 x (684x10^-9) x (0.49)]/(0.0028) = 2.394*10^-4m